OpenStudy (anonymous):

What is the diameter of a circle that has an area of 452.16 mm 2?

OpenStudy (jhannybean):

Formula for Area of a circle = \(A = \pi \cdot D\) or \(A=\pi r^2\) Given \(A= 452.16 \ mm^2\) plug A into the formula to solve for D \[\large 452.16= \pi D\]\[\large D = \dfrac{452.16}{\pi}\]

OpenStudy (anonymous):

144

OpenStudy (jhannybean):

Not exactly :\ \[\large \cfrac{452.16}{\pi}\]\[\cfrac{228700}{1589}\] and reducing that you'll find your answer :)

OpenStudy (unklerhaukus):

\[A_\text{circle}=\pi r^2=\pi (D/2)^2=\pi D^2/4\]